OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

@mulline ? @MATTW20 "?

hartnn (hartnn):

can you find slope of y = -x + 6. ?

hartnn (hartnn):

compare with y=mx+c slope =m

OpenStudy (anonymous):

-x right?

hartnn (hartnn):

slope is a number, the co-efficient of x.

hartnn (hartnn):

like for 10x, the slope will be 10

hartnn (hartnn):

so, for -x , slope =...?

OpenStudy (anonymous):

-1?

hartnn (hartnn):

yes, now you know what will be the slope of perpendicular line ?

OpenStudy (anonymous):

no...?

hartnn (hartnn):

the slope of perpendicular line is negative reciprocal( inverse.) like if the slope is 4, the slope of perpendicular line is -1/4

hartnn (hartnn):

what if the slope of line is -1 ?

OpenStudy (anonymous):

Oh! So, it is 1?

hartnn (hartnn):

yes, correct. now you have slope = m = 1 and point (x1,y1) = (-4,-2) can you find the equation?

hartnn (hartnn):

y-y1 = m (x-x1)

OpenStudy (anonymous):

the inverse of that equation?

hartnn (hartnn):

no, plug in m=1 x1 =-4, y1 =-2 in y-y1= m (x-x1)

hartnn (hartnn):

that is point slope form of equation of line

OpenStudy (anonymous):

I get it now. Thanks again. ^_^

hartnn (hartnn):

welcome ^_^ you can post what you got if you want me to verify it :)